OpenStudy (anonymous):

10. Find the perimeter of triangle CDE. Round to the nearest tenth. (1 point) 90.0 18.0 14.3 10.5

OpenStudy (anonymous):

how about a diagram?

OpenStudy (anonymous):

can u help me with it and a few more problems

OpenStudy (anonymous):

yes. you need to know how to fing the distance of a line. do you know the distance formula?

OpenStudy (anonymous):

um no??

OpenStudy (anonymous):

ok\[d=\sqrt{(x-x1)^{2}+(y-y1)^{2}}\]

OpenStudy (anonymous):

it's from the Pythagorean theorem

OpenStudy (anonymous):

ok can you please tell me the answer??

OpenStudy (anonymous):

I would appreciate it soooo much!!

OpenStudy (anonymous):

cd=5

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

the next side is \[\sqrt{45}=\sqrt{9}\sqrt{5}=3\sqrt{5}\]

OpenStudy (anonymous):

ok and the last side

OpenStudy (anonymous):

the last one is \[\sqrt{40}=\sqrt{4}\sqrt{10}=2\sqrt{10}\] so add them all up

OpenStudy (anonymous):

and the answer will be ????.....

OpenStudy (anonymous):

i did not pull out a calculator but it has to be 18

OpenStudy (anonymous):

thank so much can you help me with a few more

OpenStudy (anonymous):

yes but you must do some of the calculations

OpenStudy (anonymous):

trust me i am doing almost all of the problems I just got stuck on a few

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

17. The length of the hypotenuse of a 30-60-90 triangle is 28 m. Find the length of the side opposite the 30 angle. (1 point) 7 m 14 m 14 sqrt 3m 14 sqrt 2 m

OpenStudy (anonymous):

14

OpenStudy (anonymous):

14 m

OpenStudy (anonymous):

r u sure

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

19. Use the diagram to find sin X as a fraction in simplest form. 3/4 4/5 3/5 4/3

OpenStudy (anonymous):

hello r u still there can u please answer it fast i rneed it now PLEASE I am /begging u

OpenStudy (anonymous):

48/60=4/5

OpenStudy (anonymous):

are you sure it is not 3/5

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

22. Find the value of a in the diagram of the right triangle. Round to the nearest tenth.

OpenStudy (anonymous):

21.6 24.2 5.0 12.3

OpenStudy (anonymous):

hello r u still there can u please answer it fast i rneed it now PLEASE I am /begging u

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

24.2

OpenStudy (anonymous):

thank u soo much

OpenStudy (anonymous):

27. Zack is building a model sailboat. The mast will have two inches of height below the base of the main sail, as shown in the diagram below. He wants the base of the sail to have a length of 11 inches and angle s to measure 49. a. Write an equation that can be used to find the length of the side of the triangular sail opposite angle s . b. Find the length of the side of the triangle opposite angle s, to the nearest tenth. c. What will be the height of the mast, to the nearest tenth?

OpenStudy (anonymous):

can u please answer this on quickly 2. I would so appreciate it. this one and one more and thats it. U have helped me more than enough!

OpenStudy (anonymous):

length=tan49*11 12.65 14.7

OpenStudy (anonymous):

are the answers in abc order?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok last one and I owe u BIG time thank you so much!! :)

OpenStudy (anonymous):

28. The City Commission wants to construct a new street that connects Main Street and North Boulevard, as shown in the diagram below. a. What will be the length of the new street? Explain how you found this length. b. The construction cost has been estimated at $140 per linear foot. Estimate the cost for constructing the street. Explain your method for finding the estimate.

OpenStudy (anonymous):

a--use the Pythagorean theorem 9^2+4^2=81+16=97= of the length of the new street squared so length of new street is sqrt 97 b--140* sqrt97

OpenStudy (anonymous):

would b = 1378

OpenStudy (anonymous):

yes